OpenStudy (anonymous):

The school has 800 students with 20 students on the gymnastic team and 10 students on the chess team (including 3 students who are on both teams). How many students in the school are not members of either the gymnastic team or the chess team? {767} {770} {773} {776}

OpenStudy (anonymous):

So we have the number 800, 20 and 10. BUT there are three students who are in the same teams, meaning, we can't count these students twice So we have to subtract 3 from either 10 or 20 Let's chose 20 20 - 3 = 17 So: 17 + 10 = 27 800 - 27 = 773

OpenStudy (anonymous):

thank you so very much

OpenStudy (anonymous):

yw :)